# Emergent Hierarchical Reasoning in LLMs through Reinforcement Learning

### v2

@@ 4 Experiments
To validate our hypothesis that focusing optimization on the strategic learning bottleneck enhances reasoning, we conduct a comprehensive set of experiments. We evaluate our proposed method, Hierarchy-Aware Credit Assignment (HICRA), against strong baselines on a suite of complex reasoning benchmarks, followed by in-depth analyses to dissect the underlying learning dynamics.
4.1 Experimental Setup
Models
Our experiments are conducted on four open-source model families, including Qwen2.5-7B (Yang et al., 2024), Qwen3-4B (Yang et al., 2025), LLama-3.1-8B (Grattafiori et al., 2024), Qwen2.5-VL-7b (Yang et al., 2024) and MiMO-VL-7B (Xiaomi, 2025). We use both base and instruction-tuned variants to ensure our findings are consistent and test generalization to vision-language models.
Training Datasets and Benchmarks
[…]
Baselines and Implementation.
We compare HICRA against three primary baselines: the Base model (before RL), the widely adopted GRPO baseline with clip-higher (Yu et al., 2025) by default, and Entropy Regularization: GRPO with an additional regularization loss on token-level entropy (Cheng et al., 2025). For HICRA, we set the amplification hyperparameter $\alpha$ to $0.2$ and identify planning tokens using the Strategic Grams (SGs) methodology detailed in Section 2.1. For all experiments, we increase the training context length from 16K to 32K when the response clip rate exceeds 20% (Luo et al., 2025). For the specific experiments on Llama-3.1-Instruct, we add a dynamic filtering mechanism (Yu et al., 2025) based on GRPO Clip-Higher due to significant vanishing advantanges (Wang et al., 2025c).
4.2 Main Results
@@
Our primary results, summarized in Table 1, Table 2, show that HICRA consistently and outperforms both the GRPO baselines across text-only models and vision-language models on various benchmarks. On the strongest base model, Qwen3-4B-Instruct, HICRA’s gains demonstrate that even on highly capable models, selectively amplifying the learning signal for strategic reasoning yields substantial improvements. This trend holds for non-instruct-tuned models as well, providing strong empirical evidence for our central claim: by identifying and focusing on the emergent strategic bottleneck, HICRA accelerates the development of advanced reasoning abilities more efficiently than agnostic methods.
@@ 4.3 Analysis of RL’s Impact on Reasoning
To validate our hypothesis and the design of HICRA, we conduct a series of analyses to dissect how RL improves reasoning. Our analysis follows a clear logical thread: first, we have linked strategic planning to reasoning through analyses of the training dynamics in Section 2.2; second, we verify the key effects of RL by showing the frequency of different errors throughout training (Section 4.3.1); we then justify the effectiveness of HICRA in exploration by comparing with standard entropy-regularzed baselines. Finally, we examine the limitations of our approach.
@@ 4.3.1 Mastery of Strategic Planning Unlocks Improved Reasoning during RL
Figure 5: HICRA (red) consistently achieves higher semantic entropy than the GRPO baseline (gray), indicating more diverse strategic exploration.
@@
4.3.3 HICRA’s Limitation: Dependency on a Procedural Foundation
Figure 7: HICRA on Llama-3.1-Instruct-8B.
4.4 Semantic Entropy: A Compass for Strategic Exploration
@@
Limitations of Token Entropy and Pass@K.
As shown in Figure 8 for MiMO-VL-7B, token-level entropy “collapses” for both HICRA and GRPO, simply because the vast majority of low-level tokens are doomed to become certain, thus pulling the average token entropy down. However, this decrease in token entropy might mislead researchers to suggest that exploration has ceased. Similarly, the Pass@8 (Training) metric quickly saturates, rendering it useless for distinguishing the ongoing learning dynamics.
Semantic Entropy as the Differentiator.
In the same experiment, semantic entropy tells a more accurate story. It remains high, indicating continued exploration of diverse reasoning strategies. Crucially, HICRA consistently maintains a higher semantic entropy than GRPO, and this advantage directly correlates with its superior final validation accuracy. This also demonstrates the generality of our approach, extending effectively to multimodal reasoning tasks on vision-language models like MiMO-VL-7B.
4.5 Planning Tokens vs. High-Entropy ”Fork” Tokens
@@
Figure 9 reveals a crucial asymmetry. While a majority of planning tokens exhibit high entropy (aligning with their role as points of strategic choice), the reverse is not true: most high-entropy tokens are not planning tokens. This finding highlights the limitations of using high entropy as a standalone proxy for strategic function. High token-level entropy ensures sampling diversity, but it does not guarantee semantic function. Many high-entropy tokens may correspond to variations in phrasing or calculation that do not alter the high-level reasoning path. In contrast, our approach identifies tokens based on their functional role in orchestrating the solution, providing a more direct and reliable signal for strategic credit assignment.
5 Related Works
@@
6 Conclusion and Promising Directions
This work tackles the opacity of how Reinforcement Learning enhances the complex reasoning abilities of LLMs. We move beyond treating the learning process as a monolith, uncovering a coherent, two-phase dynamic: an initial consolidation of low-level procedural skills followed by a decisive shift where performance gains are driven by the exploration and mastery of high-level strategic planning. This emergent reasoning hierarchy provides a unified framework that explains puzzling phenomena like “aha moments,” “length-scaling,” and complex entropy dynamics.
Our core insight – that a strategic bottleneck is the primary driver for advanced reasoning – led to the development of HICRA (Hierarchy-Aware Credit Assignment). By concentrating optimization pressure on functionally-defined planning tokens, HICRA significantly outperforms strong baselines like GRPO, demonstrating that targeted credit assignment is more efficient than agnostic approaches. Furthermore, we validate semantic entropy as a far more reliable compass for measuring strategic exploration than misleading aggregate metrics like token-level entropy.
Our findings open several promising avenues for future research, suggesting a paradigm shift in how we approach RL for language models.
Rethinking Action in RL for LMs: Our analysis indicates that the token, as the fundamental unit of action, is too granular and often lacks semantic weight. A single token rarely determines the subsequent logical direction of a complex thought process. Future work should explore a higher-level action space, where an ”action” could be a semantic concept, a reasoning step (e.g., ’apply theorem X’), or a complete phrasal unit (n-gram). This would align the optimization process more closely with the cognitive steps of reasoning and could lead to more stable and efficient learning.
Dynamically Adaptive Hierarchical Methods: HICRA focused learning efforts on the strategic bottleneck. A more powerful approach should be capable of diagnosing the current learning bottleneck, which can adaptively shift optimization pressure between procedural execution and strategic planning based on real-time diagnostics of the model’s weaknesses, perhaps through a curriculum learning framework or meta-learning.
Finer-Grained, Process-Oriented Rewards: This work justifies a move away from sparse, outcome-based rewards (correct vs. incorrect final answer) towards process-oriented reward models. Such models could provide fine-grained credit assignment by, for example, rewarding a brilliant strategic choice even if it is marred by a minor calculation error. This decouples the evaluation of planning from execution, allowing the model to learn both skills more effectively.
Generalizing the Reasoning Hierarchy Beyond Mathematics: The functional hierarchy between planning and execution is likely a universal feature of complex, multi-step generation. Investigating whether similar dynamics emerge in domains like code generation , agentic tool-use, long-form generation, etc, could yield novel, domain-specific optimization strategies.
Appendix A Full Training Dynamics
@@
  • For some models, the initial skill-consolidation phase is brief or absent. In the cases of the Vision-Language Models (Qwen2.5 VL-Instruct and MiMO VL-Instruct) and Qwen3 4B-Instruct, the exploration of strategic planning begins almost immediately at the start of training. This is evidenced by a significant and immediate rise in the semantic entropy of strategic grams, which occurs in tandem with a rapid boost in validation accuracy. We conjecture this is because: (a) for the VL scenarios, publicly available datasets is learned quickly by state-of-the-art models (Wang et al., 2025c); (b) strong base models like Qwen3 4B-Instruct already possess a solid foundation of low-level skills and primarily need to adapt to formatting before focusing on higher-level strategic planning.
@@
  • Token-level entropy is misleading for policy exploration. This observation holds true across all experiments. The Qwen3 4B-Instruct model offers a particularly stark example: its token-level entropy remains almost perfectly flat, while its semantic entropy (diversity of strategic grams) consistently increases throughout training. This contrast highlights that the variety of semantic structures a model learns is completely different from the statistical uncertainty of its next-token predictions.
    The core difference is about scale: token-level entropy measures the uncertainty of every next-token, including low-level tokens such as formatting, executions that are doomed to become confident throught training. In contrast, semantic entropy measures the diversity of the overall ideas being expressed. A model can be very predictable in its next-token choice under a given context but still create a wide variety of different arguments or structures.
Appendix B The Distribution Matching Perspective of Policy Gradients
@@
Therefore, the PPO objective is essentially solving a distribution matching problem toward a target distribution shaped by the advantage function.

### v3

@@ 4 Experiments
Models and Datasets. Our experiments use open-source models including Qwen2.5-7B (Yang et al., 2024), Qwen3-4B (Yang et al., 2025), LLama-3.1-8B (Grattafiori et al., 2024), and VLMs like Qwen2.5-VL-7b (Yang et al., 2024) and MiMO-VL-7B (Xiaomi, 2025), covering both base and instruction-tuned variants. We train on established reasoning datasets DAPO (Yu et al., 2025), DeepScaleR (Luo et al., 2025) and ViRL39K (Wang et al., 2025c) for VLMs.
@@
Benchmarks and Baselines. We evaluate on a suite of challenging text-only (e.g., AIME24, AIME25 (Mathematical Association of America, 2024), Math500 (Lightman et al., 2023), AMC23, Minerva (Lewkowycz et al., 2022), and Olympiad (He et al., 2024)) and multimodal (e.g., MathVista (Lu et al., 2023), MathVerse (Zhang et al., 2024), MathVision (Wang et al., 2024), EMMA (Hao et al., 2025)) benchmarks. We adopt the evaluation protocols of Deepseek R1, using Pass@1 with random samplings. We compare HICRA against three primary baselines: the Base model (before RL), the widely adopted GRPO baseline with clip-higher (Yu et al., 2025) by default, and Entropy Regularization: GRPO with an additional regularization loss on token-level entropy (Cheng et al., 2025). A comprehensive description of our evaluation protocol, training implementation, and additional model-specific details can be found in the Appendix.
4.1 Main Results
@@
Our primary results, summarized in Table 1 and Table 2, show that HICRA consistently and outperforms both the GRPO baselines across text-only models and vision-language models on various benchmarks. On the strongest base model, Qwen3-4B-Instruct, HICRA’s gains demonstrate that even on highly capable models, selectively amplifying the learning signal for strategic reasoning yields substantial improvements. This trend holds for non-instruct-tuned models as well, providing strong empirical evidence for our central claim: by identifying and focusing on the emergent strategic bottleneck, HICRA accelerates the development of advanced reasoning abilities more efficiently than agnostic methods.
@@ 4.2 Analysis of RL’s Impact on Reasoning
We conduct a series of analyses to dissect how RL improves reasoning. First, we have linked strategic planning to reasoning through analyses of the training dynamics in Section 2.2; Second, we verify the key effects of RL by showing the frequency dynamics of different errors throughout training (Section 4.2.1); we then justify the effectiveness of HICRA in exploration by comparing with standard entropy-regularized baselines. Finally, we compare different ways of tracking exploration for RL performance and identifying critical tokens for RL.
@@ 4.2.1 Mastery of Strategic Planning Unlocks Improved Reasoning during RL
Figure 6: HICRA improves GRPO Clip-Higher via more diverse strategic exploration.
@@ 4.2.3 Semantic Entropy: A Compass for Strategic Exploration
Limitations of Token Entropy and Pass@K. As shown in Figure 8 for MiMO-VL-7B, token-level entropy “collapses” for both HICRA and GRPO, simply because the vast majority of low-level tokens are doomed to become certain, thus pulling the average token entropy down. However, this decrease in token entropy might mislead researchers to suggest that exploration has ceased. Similarly, the Pass@8 (Training) metric quickly saturates, rendering it useless for distinguishing the ongoing learning dynamics.
Semantic Entropy as the Differentiator. In the same experiment, semantic entropy tells a more accurate story. It remains high, indicating continued exploration of diverse reasoning strategies. Crucially, HICRA consistently maintains a higher semantic entropy than GRPO, and this advantage directly correlates with its superior final validation accuracy. This also demonstrates the generality of our approach, extending effectively to multimodal reasoning tasks on vision-language models like MiMO-VL-7B.
4.2.4 Planning Tokens vs. High-Entropy ”Fork” Tokens
Figure 10 and Figure 9 reveal a crucial asymmetry. While a majority of planning tokens exhibit high entropy (aligning with their role as points of strategic choice), the reverse is not true: most high-entropy tokens are not planning tokens. This finding highlights the limitations of using high entropy as a standalone proxy for strategic function. High token-level entropy ensures sampling diversity, but it does not guarantee semantic function. Many high-entropy tokens may correspond to variations in phrasing or calculation that do not alter the high-level reasoning path. In contrast, our approach identifies tokens based on their functional role in orchestrating the solution, providing a more direct and reliable signal for strategic credit assignment.
@@
5 Related Work
@@
6 Conclusions
Our work establishes that reinforcement learning uncovers an emergent functional reasoning hierarchy in language models, demonstrating the critical performance bottleneck shifting from procedural skill to strategic exploration. This insight leads to our approach, HICRA, which demonstrates that specialized credit assignment targeting this strategic bottleneck yields more effective training. Extensive experiments validate the effectiveness of HICRA and offer deep insights into advanced reasoning through strategic exploration.
Our work opens several future research directions. First, it suggests a paradigm shift away from treating all tokens equally and prompts a rethinking of the action space away from individual tokens toward semantic, strategic units. Second, it calls for developing process-oriented approaches capable of valuing correct strategic choice even if the final answer is flawed. Finally, the likely universality of this reasoning hierarchy in complex reasoning tasks suggests that applying these principles to domains like code generation and agentic tool-use is a valuable path forward.
Appendix A SG Construction
This pipeline result in the following collection of SGs.
Listing 1: Strategic Grams
1{’## step’, ’a good starting point is’, ’a more direct approach is’, ’a more straightforward approach’,
2’a simpler approach is’, ’alright’, ’alternatively’, ’an alternative path is’, ’an error in the thought process’,
3’analyze the’, ’analyzing the given’, ’and then find’, ’another approach’, ’are looking for’,
4’based on the given’, ’break down the problem’, ’break it down’, ’break it down into manageable steps’,
5’but’, ’but wait’, ’but why’, ’but without more information’, ’can be rewritten as’,
6’can conclude that’, ’can see that’, ’check if’, ’consider the case where’, ’consider the properties of’,
7’correct the approach’, ’define the variables’, ’denote’, ’determine how many’, ’directly address the problem’,
8’does this hold true?’, ’does this make sense?’, ’double-checking the logic’, ’finally need to’,
9’finally we need to’, ’find a simpler’, ’find a way to’, ’find out how many’, ’find the critical points’,
10’first need to’, ’follow these steps’, ’for simplicity’, ’from earlier we have’, ’from the above’,
11’from this, it follows that’, ’from this, we can infer’, ’given the complexity’, ’given the complexity of’,
12’given the constraints’, ’given the nature of’, ’go back to the’, ’goal is to’, ’hmm,’, ’hold on’,
13’however’, ’however, we need to’, ’i might have made an error’, ’i need to re-evaluate’, ’i should verify this result’,
14’identify the’, ’identify the given information’, ’if that doesn\’t work, we can’, ’if we consider’,
15’in a way that’, ’in the context of’, ’is there a simpler method?’, ’is there a simpler way?’,
16’it logically follows that’, ’it seems’, ’it\’s better’, ’let me’, ’let me pause and think’,
17’let me rethink this’, ’let me verify’, ’let\’s’, ’let\’s analyze the possibilities’, ’let\’s assume’,
18’let\’s backtrack’, ’let\’s break this down’, ’let\’s check our work’, ’let\’s check the constraints again’,
19’let\’s consider another case’, ’let\’s denote’, ’let\’s double-check’, ’let\’s explore a different possibility’,
20’let\’s formulate a plan’, ’let\’s go back a step’, ’let\’s outline the steps’, ’let\’s pause and think’,
21’let\’s reconsider’, ’let\’s try a different angle’, ’let\’s validate this’, ’looking back at the’, ’maybe’,
22’maybe i can’, ’my previous step was flawed’, ’need to’, ’need to account for’, ’need to analyze’,
23’need to check’, ’need to consider’, ’need to count’, ’need to determine’, ’need to ensure’, ’need to express’,
24’need to find’, ’need to follow’, ’need to identify’, ’need to minimize’, ’need to reconsider’, ’need to show’,
25’need to solve’, ’need to think about’, ’need to understand’, ’need to use’, ’next’, ’note that’, ’now’,
26’now let’, ’now need to’, ’now we need to’, ’okay’, ’on second thought’, ’on the other hand’,
27’one way to’, ’our strategy is’, ’perhaps’, ’perhaps i can’, ’problem is asking’, ’problem states that’,
28’proceed with the following’, ’rearrange the equation’, ’recall that’, ’referring to a previous step’,
29’revisiting the initial assumption’, ’rewrite the equation’, ’says that’, ’seems a bit complicated’,
30’should consider’, ’should focus on’, ’should look for’, ’similarly’, ’simplify the problem’, ’since’,
31’so’, ’so after’, ’so again’, ’so the question becomes’, ’so, yes’, ’something is wrong here’, ’specifically’,
32’start by’, ’states that’, ’step by step’, ’step by step reasoning’, ’step by step solution’,
33’step-by-step reasoning’, ’that can\’t be right’, ’that seems’, ’that was a mistake’,
34’that assumption was incorrect’, ’the correct approach is’, ’the core idea is’, ’the first step is’,
35’the key insight is’, ’the key is to realize’, ’the key to solving this’, ’the logical flow is’,
36’the next step is’, ’the path to the solution’, ’the plan is to’, ’the problem asks for’,
37’the problem is about’, ’the problem mentions’, ’the problem says’, ’the problem states’,
38’there is a mistake’, ’there seems to be’, ’therefore’, ’think of this as’, ’this allows us to’,
39’this approach isn\’t working’, ’this approach seems’, ’this can be seen as’, ’this implies’,
40’this implies that’, ’this is because’, ’this is not the correct approach’, ’this isn\’t leading anywhere’,
41’this leads to’, ’this leads us to’, ’this logically leads to’, ’this means’, ’this means that’,
42’this seems a bit’, ’this suggests’, ’this suggests a path’, ’this suggests that’, ’thus’, ’to confirm’,
43’to consider the constraints’, ’to determine’, ’to do this’, ’to ensure’, ’to ensure correctness’,
44’to find’, ’to make it easier’, ’to proceed’, ’to see if’, ’to solve this problem’, ’to verify’,
45’try to’, ’understand the given information’, ’understanding the problem’, ’understanding the problem first’,
46’upon closer inspection’, ’use the concept of’, ’use the fact’, ’use the fact that’, ’use the method of’,
47’use the properties of’, ’verify the solution’, ’wait’, ’wait, but’, ’wait, no’, ’wait, that\’s not right’,
48’want to find’, ’we are dealing with’, ’we can’, ’we can approach this’, ’we can conclude’, ’we can deduce’,
49’we can infer’, ’we can see’, ’we can start by’, ’we can think of this as’, ’we can use’, ’we know’,
50’what am i missing?’, ’what happens if’, ’what if we assume’, ’what if we try’, ’what is being asked’,
51’which means’, ’will consider the’, ’work our way’}
(a) Full SGs
(b) 30% SGs randomly dropped
Figure 11: Sensitivity Analysis of randomly dropping 30% strategic grams on Qwen3-4B-Base training dynamics. The semantic entropy curve remain identical.
Appendix B Sensitivity Analysis of SGs
(a) Full SGs
(b) 30% SGs randomly dropped
Figure 12: Sensitivity Analysis of randomly dropping 30% strategic grams on Qwen2.5-7B-Base training dynamics. The semantic entropy curve remain identical.
This automated procedure is designed to yield a high-precision functional proxy for strategic planning, not an exhaustive lexicon of all possible SGs. We set reasonable hyper-parameters for identifying SGs, and we contend that the resulting SG collection is sufficiently representative to reveal the core learning dynamics. To validate this claim, we conduct a sensitivity analysis by randomly removing 30% of the identified SGs and re-running our main analysis. As shown in Figure 11 and Figure 12, the semantic entropy curves remain qualitatively identical, and the curves for perplexity and token entropy only slightly change. Semantic entropy here calculates the entropy of frequency distribution, which itself is not sensitive to slight changes in the frequency mass, as long as there are sufficient numbers of bins. This demonstrates the robustness of our SG identification and the findings derived from it.
@@ Appendix C Full Training Dynamics
  • The initial skill-consolidation phase might be brief or absent for some models. In the cases of the Vision-Language Models (Qwen2.5 VL-Instruct and MiMO VL-Instruct), Qwen3 4B-Instruct, Deepseek-Distill-Llama-8B, the exploration of strategic planning begins almost immediately at the start of training. This is evidenced by a significant and immediate rise in the semantic entropy of strategic grams, which occurs in tandem with a rapid boost in validation accuracy. We conjecture this is because: (a) for the VL scenarios, publicly available datasets is learned quickly by state-of-the-art models (Wang et al., 2025c); (b) strong base models like Qwen3 4B-Instruct already possess a solid foundation of low-level skills and primarily need to adapt to formatting before focusing on higher-level strategic planning.
@@
  • Token-level entropy is misleading for policy exploration. This observation holds true across all experiments. The Qwen3 4B-Instruct model offers a particularly stark example: its token-level entropy remains almost perfectly flat, while its semantic entropy (diversity of strategic grams) consistently increases throughout training. This contrast highlights that the variety of semantic structures a model learns is completely different from the statistical uncertainty of its next-token predictions. Figure 4 illustrates the differences of the two entropy.
    The core difference is about scale: token-level entropy measures the uncertainty of every next-token, including the vast amount of low-level tokens such as formatting, executions that are doomed to become confident throughout training. In contrast, semantic entropy measures the diversity of the overall meanings being expressed. A model can be very predictable in its next-token choice under a given context but still create a wide variety of different arguments or structures.
  • Lack of Strategic Exploration Hinders Sustained Improvement in Llama Models. We observe that the Llama-3.1-8B-Base model initially focuses almost exclusively on consolidating low-level procedural skills, a phase marked by decreasing perplexity and token entropy on execution tokens. However, once the performance gains from this procedural refinement diminish, the model fails to pivot towards exploring high-level planning strategies. This leads to performance stagnation and, eventually, degradation.
    This behavior stands in stark contrast to the more successful Deepseek Distilled Llama model, which engages in high-level strategic exploration from the very beginning of training, bypassing a distinct procedural consolidation phase. We hypothesize that for the standard Llama models, the intense initial focus on procedural correctness prematurely collapses the diversity of high-level reasoning strategies. By the time low-level skills are mastered, the model has likely converged on simpler reasoning patterns, which inhibits its ability to subsequently discover and adopt more complex and effective problem-solving approaches.
Appendix D The Distribution Matching Perspective of Policy Gradients
@@
Therefore, the PPO objective is essentially solving a distribution matching problem toward a target distribution shaped by the advantage function. It follows that a standard policy gradient (Williams, 2004) update nudges the policy $\pi_{\theta_{old}}$ toward an implicit target distribution $\pi^{\ast}$ defined by the advantage function. After the gradient update, the target distribution becomes the new policy sampled for exploration. Therefore, adjusting the advantage function (or credit assignment) essentially shapes the exploration policy during RL training.
Appendix E Extended Materials of Experiments
E.1 Experimental Setups
Training Datasets and Benchmarks
[…]
Baselines and Implementation.
We compare HICRA against three primary baselines: the Base model (before RL), the widely adopted GRPO baseline with clip-higher (Yu et al., 2025) by default, and Entropy Regularization: GRPO with an additional regularization loss on token-level entropy (Cheng et al., 2025). For HICRA, we set the amplification hyperparameter $\alpha$ to $0.2$ and identify planning tokens using the Strategic Grams (SGs) methodology detailed in Section 2.1. For all experiments, we increase the training context length from 16K to 32K when the response clip rate exceeds 20% (Luo et al., 2025). For the specific experiments on Llama-3.1-Instruct, we add a dynamic filtering mechanism (Yu et al., 2025) based on GRPO Clip-Higher due to significant vanishing advantanges (Wang et al., 2025c). We use two to four sets of eight A100 (80G) for training all models, and we stop the experiments if performance continues to degrade during extended training.
Prompt Used for Error Analysis. Listing LABEL:lst:error shows the prompt for identifying error types in our experiments.
Listing 2: Instruction for Identifying Error Types
1The goal of this analysis is to evaluate a student’s solution to a math problem and determine the key reason for any failure. We will categorize failures into two main types: those stemming from high-level strategic/conceptual errors and those from low-level execution/concrete mistakes.
2
3### Instruction
4
51. **Understand the Problem:** Read the math problem carefully to grasp what is being asked.
62. **Analyze the Reference Solution:** Review the correct solution step-by-step. Identify the core concepts and high-level strategies used to arrive at the correct answer. Note the key principles or formulas applied.
73. **Analyze the Student’s Solution:** Read the student’s solution from start to finish.
8 * Compare the student’s approach to the reference solution. Where did they diverge?
9 * Look for specific errors. Are they mistakes in calculation (e.g., \$2+2=5\$)? Are they mistakes in variable substitution (e.g., using the wrong value for \$x\$)? ...
10 * Determine the **single most significant reason** the student’s final answer is incorrect.
11
124. **Determine the Failure Type:** Based on your analysis, choose one of the following four options.
13 * **(A) Failure due to dummy reasons:** The student solution does not seem related with the problem, or the student lacks a basic understanding about the problem, resulting in an invalid solution.
14 * **(B) Failure in Low-Level Executions and Procedural Fidelity:**
15 - the error occurs on a **concrete/specific** step,
16 - the errors relate with a failure of performing basic math skills
17 - the errors mistake at executions or procedural analysis of the high-level plans/strategies
18 For example, errors in rewriting equations, calculation, doing algebra, variable substition, solving equations, procedural analysis, etc.
19 * **(C) Failure due to Flaws in High-Level Plans:** When the problem is complex, it requires the student to employ certain high-level plans, strategic moves or regular self-critiques. The student had conceptual errors, used an incorrect high-level approach or mistake at certain strategic moves to solve the problem.
20 * **(D) The student answer is indeed acceptable:** The student’s solution is correct, even if it differs from the reference solution’s method.
21
[…]
23
24Here is the problem: {}
25
26Here is the reference solution: {}
27
28Here is the student solution:
29{}’
E.2 HICRA Presumes a Dependency on a Procedural Foundation
Figure 14: HICRA on Llama-3.1-Instruct-8B.